a certain amount of a metal whose equ.mass os 28 displaces 0.7 L of H2 at s.t p from an acid hence mass of the element is

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Hiii 
Vol Of H2 at 1 atp = 22.4L
Mole of 0.7 L H2 at STP = 0.7/22.4 = 0.0313 moles.
weight of element = Moles H2 * eq . weight of metal * mass of H2 
= 1.7528 gm
Arun
4 years ago

1 mole of H2 at STP occupies volume = 22.4 L

No. of moles of 0.7 L H2 at STP = 0.7/22.4 = 0.0313 moles.

Weight of element (X) = Moles of H2 x eq. wt. of the metal x mass of H2

X = 0.0313 x 28 x 2

X = 1.7528 grams.